Grade 11General Physics

A tungsten wire of length 20cm is stretched by 0.1 cm. Find the strain on the wire

Deepakkumar
9 Years agoGrade 11

1 Answer

Abhiram
9 Years ago
Given that,
Length of the wire, L=20 cm
The wire is stretched,the elongation in the wire,{\mathit{\Delta}}L=0.1 cm
Now,Strain={\mathit{\Delta}}L/L 
Strain= 0.1/20= 5×10-3